### v1

# Green-LLM: Optimal Workload Allocation for Environmentally-Aware Distributed Inference

Abstract
This letter investigates the optimal allocation of large language model (LLM) inference workloads across heterogeneous edge data centers (DCs) over time. Each DC features on-site renewable generation and faces dynamic electricity prices and spatiotemporal variability in renewable availability. The central question is: How can inference workloads be optimally distributed to the DCs to minimize energy consumption, carbon emissions, and water usage while enhancing user experience? This letter proposes a novel optimization model for LLM service providers to reduce operational costs and environmental impacts. Numerical results validate the efficacy of the proposed approach.
Index Terms:
LLM inference, sustainable AI, renewable energy, carbon and water footprint reduction, edge data centers.
I Introduction
Rapid advancements in LLMs have revolutionized natural language processing, enabling AI systems to achieve human-level performance across numerous tasks, such as text generation, sentiment analysis, and machine translation [1]. LLMs operate through two critical phases: training and inference. The training phase, typically a one-time operation, can span weeks or even months and demands massive computational resources. In contrast, inference occurs in real-time and directly affects user experience, where high-quality service and low-latency responses are essential [2].
Despite their transformative capabilities, LLMs incur substantial operational and environmental costs, particularly given the exponential growth in model sizes. The electricity consumed by AI DCs strains power grid stability, increases carbon emissions, and elevates water usage for cooling. The widespread deployment of LLM models can also influence electricity prices, potentially escalating living costs for nearby communities. These concerns underscore the need for energy-efficient strategies that mitigate the environmental costs of LLM inference and promote the responsible AI deployment.
In LLM systems, each user query consists of input tokens (i.e., text or data provided by the user) and output tokens (i.e., the generated response). The energy required for inference is strongly influenced by token size, as larger or more complex queries demand greater computational resources [3]. While general-purpose models like GPT-4 can handle diverse requests, their large size often results in high energy consumption during inference. An alternative approach is to route requests to smaller, task-specific sub-models tailored to applications such as code generation, translation, or summarization [2]. These task-specific models, designed as specialized sub-models, consume less energy and reduce the carbon and water footprint while maintaining comparable accuracy [4, 5, 6]. Furthermore, DCs often include a mix of heterogeneous hardware, such as GPUs and CPUs of varying capabilities. Smaller LLM models can utilize less powerful hardware while ensuring efficient inference performance, resulting in lower energy consumption [7].
Additionally, the LLM SP can leverage variations in electricity prices and on-site renewable energy generation across different locations to optimize its operations. Electricity prices and renewable energy outputs fluctuate significantly across both time and spatial domains. Therefore, “how to strategically allocate user workloads to DCs located in regions with abundant renewable energy and lower electricity prices.” is a key and challenging problem that needs to be addressed.
However, energy-efficient LLM inference and workload allocation are emerging topics that remain underexplored. Limited research has focused on environmentally aware LLM inference and workload allocation. [7] models the workload-dependent energy consumption and runtime of LLM inference tasks on heterogeneous GPU-CPU systems. [8] considers an energy-management framework for the LLM inference environment, which automatically reconfigures the inference cluster to optimize for energy and cost of LLM serving under the server’s performance. [9] proposes energy-efficient models that can deploy generated LLMs on resource-limited edge devices to meet the memory and timing constraints to minimize energy consumption. However, existing studies often treat networking and energy management in isolation, failing to explore the advantages of integrated systems. They largely overlook the impact of variations in electricity prices across regions and the availability of renewable energy sources. In addition, the paper may not incorporate several physical constraints related to environmental sustainability. To this end, this letter aims to tackle the economic and environmental challenges posed by large-scale LLM operations, issues that are often overlooked despite their growing significance.
Figure 1: System model
II System Model and Problem Formulation
This section formulates the sustainable workload allocation problem for distributed inference for an LLM service provider (SP). We consider a setting where the SP operates a network of edge and cloud DCs to process inference requests. In making allocation decisions, the SP must account for the complex interdependencies among user demand, resource capacity, renewable energy availability, carbon emissions, and freshwater consumption. The SP aims to develop an optimized allocation strategy that minimizes energy consumption, ensures high service quality, and promotes sustainable operations.
II-A System Model
In this model, we consider an LLM SP that manages a set of geo-distributed DCs to serve users located across various areas. Let $\mathcal{I}$ and $\mathcal{J}$ denote the sets of areas and DCs, respectively, with area index $i\in\mathcal{I}$ and DC index $j\in\mathcal{J}$. During inference, each user generates a query of type $k\in\mathcal{K}$, where different query types correspond to distinct LLM sub-models optimized for tasks, such as translation, summarization, or code generation. To efficiently route and process these queries, the LLM SP employs a pre-trained classifier to automatically categorize incoming prompts based on their content [10].
Once classified, the query is forwarded to an appropriate DC hosting the LLM sub-model $k$, and the generated response is returned to the source area. Let $\lambda_{i,k}^{t}$ denote the total number of type-$k$ queries generated in area $i$ at $t$. We introduce a variable $x_{i,j,k}^{t}\in[0,1]$ to indicate the fraction of type-$k$ queries generated in area $i$ processed by DC $j$ at time $t$. Each query type $k$ has average input and output token lengths, denoted by $h_{k}$ and $f_{k}$, respectively. The energy and water consumption for serving a query are proportional to the input and output token counts.
To process inference workloads, DCs consume computing resources such as RAM, GPU, and CPU. Let $r\in\mathcal{R}$ indicate computing resource type and $C_{j}^{r}$ represent the capacity of resource type $r$ at DC $j$. Each DC $j$ may have on-site renewable energy sources, such as wind turbines and solar panels. Let $P_{j,t}^{\sf w}$ denote the amount of renewable energy generated at DC $j$ at time $t$. The SP can also procure electricity from the main grid. To reduce both electricity cost and environmental impact, the SP should prioritize allocating inference workload to DCs with greater renewable energy availability or lower electricity prices. Given the spatial and temporal variability in electricity prices and renewable generation, workload allocation decisions must be dynamically adjusted over time.
II-B Problem formulation
In the following, we introduce cost objectives, followed by physical constraints considered in the proposed model.
1) Power procurement cost: To power the AI DCs, the SP can procure electricity from the main grid, which incurs a cost:
$\mathcal{C}_{1}({\mathbf{P^{g}}})=\sum_{j,t}c_{j}^{t}P_{j,t}^{\sf g},$ (1)
where $P_{j,t}^{\sf g}$ represents the total procured electricity from the grid and $c_{j}^{t}$ is the electricity price at DC $j$ at time $t$.
2) Carbon emission cost: Let $\theta_{j}^{t}$ denote the carbon intensity, i.e., the emission factor per unit of electricity procured from the grid at DC $j$ at time $t$. The resulting carbon emission at DC $j$ at time $t$ is $l_{j}^{t}=\theta_{j}^{t}P_{j,t}^{\sf g},\forall j,t$. Let $\delta_{j}$ represent the carbon price (e.g., via carbon tax or cap-and-trade) at location $j$. The total carbon emission cost is given by [11, 12]:
$\mathcal{C}_{2}({\mathbf{P^{g}}})=\sum_{j,t}\delta_{j}l_{j}^{t}=\sum_{j,t}\delta_{j}\theta_{j}^{t}P_{j,t}^{g}$ (2)
3) Delay penalty: We consider transmission delay, propagation delay, and processing delay. Let $\beta_{i,k}^{t}$ denote the average size of tokens (in bits) at time $t$, and $B_{i,j}$ denote the available bandwidth on the link between $i$ and $j$. The average transmission delay for query type-$k$ from area $i$ is:
$D_{i,k,t}^{\sf tran}=\sum_{j}\frac{\beta_{i,k}^{t}(h_{k}+f_{k})x_{i,j,k}^{t}}{B_{i,j}}.$ (3)
Let $d_{i,j}$ denote the network delay between $i$ and $j$. The average propagation delay for query type-$k$ from area $i$ is:
$D_{i,k,t}^{\sf prop}=\sum_{j}d_{i,j}x_{i,j,k}^{t}.$ (4)
Let $v_{j,k}$ be the average processing delay per unit token of query type-$k$ at DC $j$. The average processing delay for query type-$k$ from area $i$ is:
$D_{i,k,t}^{\sf proc}=\sum_{j}v_{j,k}(h_{k}+f_{k})\lambda_{i,k}^{t}x_{i,j,k}^{t}.$ (5)
Then, the total delay cost can be expressed as:
$\mathcal{C}_{3}({\mathbf{x}})=\sum_{i,k,t}\rho^{k}\Big{(}D_{i,k,t}^{\sf tran}+D_{i,k,t}^{\sf prop}+D_{i,k,t}^{\sf proc}\Big{)},$ (6)
where $\rho^{k}$ is the unit delay penalty. Our problem is subject to the following constraints:
1) Power balance constraints: The energy consumption for LLM inference is closely tied to the number of input and output tokens [13]. Let $\tau_{k}=(\tau_{k}^{\sf in},\tau_{k}^{\sf out})$ denote the energy consumption coefficients per input and output token, respectively, for query type $k$, which can be estimated from historical data. The total energy consumption for inference at DC $j$ at time $t$ (i.e., $P_{j,t}^{\sf c}$) is given by [13]:
$P_{j,t}^{\sf c}=\sum_{i}\big{(}\tau_{k}^{\sf in}h_{k}+\tau_{k}^{\sf out}f_{k}\big{)}\lambda_{i,k}^{t}x_{i,j,k}^{t}.$ (7)
The Power Usage Effectiveness (PUE) metric is employed to quantify the overall energy efficiency of a DC [14]. We assume that each DC exclusively handles inference workloads. Thus, the total energy consumption at DC $j$ in time $t$ is:
$P_{j,t}^{\sf d}={\mathsf{PUE}}_{j}P_{j,t}^{\sf c},~{}~{}\forall j,t.$ (8)
Besides the main grid, each DC $j$ may have an on-site renewable energy source, generating $P_{j,t}^{\sf w}$ at time $t$. We have the following energy balance constraint:
$P_{j,t}^{\sf d}=P_{j,t}^{\sf g}+P_{j,t}^{\sf w},~{}\forall j,t.$ (9)
2) Grid capacity constraints: The power procured from the grid cannot exceed the capacity $P_{j,t}^{\sf max}$ at the point of common coupling between DC $j$ and the grid, i.e., we have:
$0\leq P_{j,t}^{\sf g}\leq P_{j,t}^{\sf max},~{}\forall j,t.$ (10)
3) Water consumption constraints: Let ${\mathsf{WUE}}$ denote the direct Water Usage Effectiveness, representing the ratio of water to IT energy usage. Direct water usage is associated with server cooling systems and can be calculated using the direct WUE. Indirect water consumption arises from electricity usage and is influenced by the local Electricity-Water Intensity Factor (EWIF), which quantifies the amount of water used per unit of electricity generated. The total water consumption at DC $j$ during time $t$ is given by [15, 16]:
$W_{j,t}=\bigg{(}\frac{{\mathsf{WUE}}_{j,t}}{{\mathsf{PUE}}_{j}}+{\mathsf{EWIF}}_{j,t}\bigg{)}P_{j,t},~{}~{}\forall j,t.$ (11)
To ensure sustainable water management, the SP may impose a limit on the total water consumption [17]:
$\sum_{j,t}W_{j,t}\leq Z.$ (12)
4) Resource constraints: To serve LLM queries, the SP must deploy appropriate LLM sub-models at the DCs. The SP must also ensure sufficient computing resources to process inference workloads. Let $\alpha_{k,r}$ denote the amount of resource type $r$ required for processing each unit of type-$k$ query. We have:
$\sum_{i,k}\alpha_{k,r}(f_{k}+h_{k})\lambda_{i,k}^{t}x_{i,j,k}^{t}\leq C_{j}^{r},~{}\forall j,r,t.$ (13a)
5) Workload allocation constraints: LLM queries must be fully served by the DCs, i.e., we have:
$\sum_{j}x_{i,j,k}^{t}=1,~{}\forall i,k,t.$ (14)
6) Delay constraints: To ensure a satisfactory user experience, the SP can impose a certain threshold $\Delta_{i,k}$ on the average total delay, i.e., we have:
$D_{i,k,t}^{\sf trans}+D_{i,k,t}^{\sf prop}+D_{i,k,t}^{\sf proc}\leq\Delta_{i,k},\forall i,k,t.$ (15)
Overall, the problem for the SP can be cast as the following multi-objective mixed integer linear program (MO-MILP):
M0: $\min_{{\mathbf{P^{g}}}}\mathcal{C}_{1}({\mathbf{P^{g}}});\min_{{\mathbf{x}}}\mathcal{C}_{2}({\mathbf{P^{g}}});\min_{{\mathbf{x}}}\mathcal{C}_{3}({\mathbf{x}})$ (16a)
[…]
To solve this MO-MILP, a common method is to assign different weights to each cost component:
$\min_{{\mathbf{P^{g}}},{\mathbf{x}}}$ $~{}~{}\sigma_{e}\mathcal{C}_{1}({\mathbf{P^{g}}})+\sigma_{c}\mathcal{C}_{2}({\mathbf{P^{g}}})+\sigma_{d}\mathcal{C}_{3}({\mathbf{x}})$ (17a)
[…]
However, determining appropriate weights in a multi-objective MILP framework can be particularly challenging, as it requires balancing heterogeneous objectives such as cost, delay, and environmental impact. To circumvent the need for manual weight tuning, we adopt a lexicographic optimization solution to solve our proposed model, as shown in Algorithm 1. This method establishes a strict priority order among the objectives and solves a sequence of MILPs, each minimizing a single objective while enforcing constraints to preserve the near-optimality of all higher-priority objectives.
Algorithm 1 Lexicographic Optimization for Green-LLM
0:  Tolerance $\epsilon$; cost parameters; priority list $\mathcal{O}=[o_{1},o_{2},o_{3}]$, where $o_{\ell}\in\{C_{\text{energy}},C_{\text{carbon}},C_{\text{delay}}\}$
1:  Initialize phase_results [$\leftarrow$[\ ]← [ ]
[…]
3:  for $\ell=1$ to $3$ do
4:     $o\leftarrow o_{\ell}$Current priority objective}
[…]
6:     for each $o^{\prime}\in\{o_{1},\dots,o_{\ell-1}\}$ do
7:        Add $C_{o^{\prime}}\leq(1+\epsilon)\cdot\texttt{optimal\_values}[o^{\prime}]$
8:     end for
9:     Update optimal_values$[o]\leftarrow C_{o}$
10:     Append results to phase_results
11:  end for
12:  return  Final decisions $(\mathbf{x},\mathbf{P}_{g})$ and phase_results
III Numerical Results
III-A Simulation setup
III-A1 Network topology
We consider a network system comprising $9$ data centers (DCs) ($J=9$) and $9$ areas ($I=9$). The network topology is based on the locations of Google Cloud DCs. Network delays between any two locations are obtained using the global ping dataset [^1] https://wondernetwork.com/pings, while electricity prices are extracted from https://www.gridstatus.io/. Carbon intensity values $(\theta_{j}^{t})$ are derived from the Google Cloud dataset [^2] https://cloud.google.com/sustainability/region-carbon. Regional carbon taxes are generated by scaling a base carbon tax of \$$\$50\mathrm{/tCO}_{2}$ by a region-specific factor [18]. In the default setting, the renewable energy generation $P_{j,t}^{\sf w}$ at each DC is simulated based on wind speeds drawn from a Weibull distribution [19] with shape $k=2$ and scale $\lambda=7$, then scaled to the range $[500,1000]$ KW to reflect a typical mid-scale wind turbine output. PUE values are taken from publicly available Google Cloud statistics [^3] https://datacenters.google/efficiency/. The WUE and EWIF values are also obtained from the public sources [20]. We consider $R=4$ different types of physical resources, with their capacities ($C_{j}^{r}$) based on specifications of selected Google Cloud DCs [^4] https://cloud.google.com/compute/docs/machine-resource.
III-A2 LLM inference workload
We generate the time-varying query demand $\lambda_{i,k}^{t}$ using a base hourly demand signal that captures 24-hour usage patterns, scaled by a region-specific population multiplier to reflect spatial heterogeneity. We consider five types of task-specific queries (i.e., $K=5$): chat, text summarization, math solving, code generation, and image generation. Each task type $k$ is associated with an average number of input tokens $h_{k}=\{40,500,30,40,30\}$ and output tokens $f_{k}=\{100,250,100,500,50\}$. Each task type is weighted by a query-type popularity factor to reflect how commonly each type appears in practice (e.g., $a^{\text{pop}}_{k}=[2.5,1.5,1.3,0.8,0.6]$). To model the temporal demand fluctuations, we modify the base demand by using a “time-of-use” demand to differentiate peak and off-peak periods. Specifically, for each area and query type, demand during peak hours (2:00 pm to 8:00 pm) is drawn from a uniform interval $[900,1000]$, while the off-peak demand is randomly generated from the interval $[500,600]$.
III-A3 Other parameters
The processing delay for each type of $k$ query is generated based on the complexity of the task, i.e., $v_{j}^{k}=[10^{-3},0.002,10^{-2},0.02,0.03]$ms/token. For simplicity in describing the results, the average delay threshold ($\Delta_{i,k}$) is set to $5s,\forall i,k$. All the experiments are conducted in Python using Gurobipy [^5] https://pypi.org/project/gurobipy/ on a desktop with an Intel Core i7-11700KF CPU and $32$GB of RAM. The source code, dataset, and models are available at: https://github.com/JJmingcc/Green_LLM.
III-B Performance evaluation
We evaluate the performance of the proposed model M0 by comparing it against the following benchmark schemes:
  • M1 (Energy-Saving Model): We prioritize allocating LLM workloads to DCs with the lowest electricity prices, where $(\sigma_{e},\sigma_{c},\sigma_{d})=(1,0,0)$.
  • M2 (Carbon-Saving Model): We prioritize allocating LLM workloads to DCs with the lowest carbon emissions, where $(\sigma_{e},\sigma_{c},\sigma_{d})=(0,1,0)$.
Models M1M2 optimize decisions by focusing exclusively on individual ecological components—namely, renewable availability, and carbon intensity. We will evaluate the performance of all models in terms of total cost and carbon emissions.
  • Total cost: $\mathcal{C}^{\sf a}=\mathcal{C}_{1}({\mathbf{P^{g}}})+\mathcal{C}_{2}({\mathbf{P^{g}}})+\mathcal{C}_{3}({\mathbf{x}})$
  • Carbon emission: $\mathcal{CO}_{2}=\mathbf{\theta}^{\top}{\mathbf{P^{g}}}.$
where decisions $({\mathbf{P^{g}}},{\mathbf{x}})$ are obtained from M0 - M2.
(a) Total cost
(b) Carbon emission
(c) Total cost (t)
(d) Carbon emission (t)
Figure 2: The impacts of the carbon intensity $\theta$
1) Varying carbon intensity: Figs.2(a)-2(b) illustrates the trade-off between carbon emission and total operational costs as a function of the carbon intensity scaling factor $\Psi_{\theta}$ across these three models. A lower value of $\Psi_{\theta}$ indicates a cleaner energy mix or reduced carbon emissions in the power grid supplying the DCs. In Fig. 2(b), M0 consistently emits significantly less carbon than M1, especially at higher carbon intensities where the gap widens substantially. While M2 achieves the lowest carbon emissions due to its carbon-minimization objective, M0 stays relatively close in emission performance—demonstrating clear carbon awareness—despite not optimizing for emissions directly. On the cost side (Fig. 2(a)), M0 significantly outperforms both M1 and M2. Figs 2(c)-2(d) present the hourly breakdown of carbon emissions and cost profiles for the three models under a moderately high carbon intensity setting ($\Psi_{\theta}$ = 1.2). M1 optimizes solely for energy cost, showing volatile and spike-prone behavior in both carbon emissions and cost. It schedules heavy workloads during carbon-intensive hours (e.g., 6:00 and 18:00–21:00), resulting in sharp emission peaks and concentrated cost surges. In contrast, M2 distributes workload more evenly across low-carbon periods, effectively reducing emissions but at the expense of higher cost and limited flexibility. M0 maintains low carbon emission levels well below M1 while incurring lower costs than M2. Without explicitly reshaping workload over time, M0’s objective structure ensures smoother and more consistent temporal performance, making it well-suited for carbon-aware yet cost-conscious deployment.
(a) Delay penalty (t)
(b) Carbon emission (t)
Figure 3: The impacts of renewable energy ($P_{w}$)
2) Varying availability of renewable energy: Figs 3(a)-3(b) analyze the impact of increasing renewable penetration, represented by $\Psi_{P_{w}}$, on the performance of the three models. As $\Psi_{P_{w}}$ increases, more renewable energy becomes available, reducing reliance on the carbon-intensive main grid. Fig. 3(b) shows that M0 maintains a more stable carbon emission pattern throughout the day, avoiding the high volatility of M1 and exhibiting lower emissions than M2 during several hours. Importantly, Fig. 3(a) shows that M0 also results in the lowest overall delay penalty, especially during peak hours, which ensures the QoS of the user.
Takeaway 1: Our proposed M0 model achieves a substantial reduction in total cost compared to M2 while incurring only a modest increase in carbon emissions.
(a) Total cost: varying $\tau$
(b) Carbon emission: varying $\tau$
(c) Delay: varying $\rho$
(d) Total cost: varying $\rho$ and $\tau$
Figure 4: Varying token size $\mathbf{\tau}$ and delay penalty $\rho$
3) Impacts of other parameters: Figs. 4(a)4(b) illustrates how variations in compute token size $\tau$ and delay penalty weight $\rho$ affect carbon emissions and operational costs. Figs. 4(a)4(b), as $\Psi_{\tau}$ increases—indicating larger token size—the total carbon emission and total cost rises sharply for all models since a higher $\Psi_{\tau}$ corresponds to longer processing durations and increased computational load. M1 is most sensitive to this growth of token size, while M2 maintains low emissions at the expense of higher cost. M0, in contrast, achieves a favorable balance—keeping emissions lower than M1 and costs lower than M2 throughout. Fig 4(c) explores the effect of increasing delay penalties ($\Psi_{\rho}$) on total cost: as expected, higher penalty weights drive up the cost across all models. Yet, M0 remains the most cost-efficient, consistently outperforming both M1 and M2. Fig.4(d) further reveals the cost sensitivity to token-related delay increases sharply when $\Psi_{\rho}$ is high. This highlights the compunding impact of long processing time and strict delay penalties on operational costs.
Takeaway 2: M0’s robustness in managing inference workloads where large token sizes and delay penalties coexist
4) Impacts of weight parameters: The comparative analysis of lexicographic and weight-vector formulations for Green-LLM reveals that the lexicographic scheme delivers strict compliance with a chosen hierarchy of objectives but induces pronounced trade-off discontinuities—total cost, carbon, or latency can swing by well over $100\%$ when the top priority shifts—whereas the weighted formulation yields much smoother response surfaces, keeping total cost within a ±$\pm 0.5\%$ band while allowing large proportional gains in the targeted metric as shown in TABLE.I and II. In practical terms, modest increases in the carbon weight (e.g., $\sigma_{c}=0.6$) halve carbon emissions for $<1\%$ extra cost, and analogous adjustments in the delay weight produce latency savings at similarly negligible budget impact, outcomes that would demand far more expensive concessions under lexicographic delay- or carbon-first orders.
Takeaway 3: Weighted model provides a Pareto-efficient, easily tunable policy lever, while lexicographic optimization should be reserved for scenarios with stringent SLA mandates that dictate hard caps on a single objective.
Priority
Order
Total
Cost
Energy
cost
Carbon
emission
Delay
penalty
E $\succ$ D $\succ$ C 394.46 193.94 8.14 192.19
E $\succ$ C $\succ$ D 395.66 194.13 7.84 193.68
D $\succ$ E $\succ$ C 628.24 425.72 22.51 180.00
D $\succ$ C $\succ$ E 642.20 443.58 18.62 180.00
C $\succ$ E $\succ$ E 404.72 207.13 2.09 195.49
C $\succ$ D $\succ$ E 404.89 207.11 2.09 195.69
TABLE I: Different priority orders in lexicographic method
Weight Vector
($\sigma_{e},\sigma_{c},\sigma_{d}$)
Total
cost
Energy
cost
Carbon
cost
Delay
penalty
(0.33, 0.33, 0.33) 392.08 197.63 5.46 189.00
(0.60, 0.20, 0.20) 394.01 194;.32 8.01 191.68
(0.20, 0.60, 0.20) 393.89 202.86 2.96 188.08
(0.20, 0.20, 0.60) 393.41 201.64 4.55 187.23
(0.45, 0.45, 0.10) 393.42 195.33 6.29 191.80
(0.45, 0.10, 0.45) 393.13 195.64 7.64 189.85
TABLE II: Different weights parameters in weighting method
IV Conclusion
This letter proposes a new optimization model for LLM SPs to efficiently allocate inference workloads across distributed, heterogeneous edge DCs. By incorporating dynamic factors such as renewable energy availability, carbon intensity, electricity prices, and water usage, the model balances environmental sustainability with operational efficiency. The results demonstrate that it reduces costs and environmental impact without compromising service quality, providing a promising direction toward sustainable AI infrastructure development.